can someone help pretty please
find the least common denominator.
17. 2m+3/m^2-25 and 4m-2/m^2+3m-10

Answer by ewatrrr(24785)


Hi
(2m+3)/(m^2-25) and (4m-2)/(m^2+3m-10)
m^2 -25 = (m+5)(m-5)
m^2+3m-10 = (m+5)(m-2)
LCD would be (m+5)(m-5)(m-2)